OpenStudy (anonymous):

In SF6, the mass of fluorine per gram of sulfur is 3.55 g F/g S. In the other compound, SFX, the mass of fluorine per gram of sulfur is 1.18 g F/g S. What is the value of X for the second compound?

OpenStudy (julian.pitalua):

\[X _{F}=(3.55gF \times32,06gS)/18.99g(F)= 6 FIRST\] Then \[X _{F}(1,18gf \times 32,06gS)/18.99= 2\]